OpenStudy (anonymous):

Simplify completely: log 8(2)

OpenStudy (anonymous):

\[\log_8(2)\]?

OpenStudy (anonymous):

solve \[8^x=2\] for \(x\)

OpenStudy (anonymous):

hint: two is the cubed root of \(8\)

OpenStudy (anonymous):

hint hint the cubed root is written in exponential form as \[\sqrt[3]{8}=8^{\frac{1}{3}}\]